A particle is moving in a straight line with initial velocity u and uniform acceleration f if sum of distances travelled in t th and t+1 the second is 100 find velocity after t th second.

Grade:11

1 Answers

Vikas TU
7 years ago
For th second,
x1 = u + f/2(2t-1).................................(1)
x2 = u +f/2(2(t+1)-1)........................(2)
x1 + x2 = 100........................(3)
soolve these three euations. and get the vlue in th  second.